# D2-Mamba: Dual-Scale Fusion and Dual-Path Scanning with SSMs for Shadow Removal

### v2

@@ Introduction
Shadows are common in natural images and often lead to visually distracting artifacts that adversely affect both human perception (Cucchiara et al. 2003; Nadimi and Bhanu 2004) and the performance of high-level vision tasks (Girshick et al. 2014; Yang et al. 2018; Özdemir and Akagündüz 2024). The goal of shadow removal is to restore clean, shadow-free images from inputs that are only partially degraded. Unlike other low-level restoration tasks, such as deblurring (Zou et al. 2021; Cho et al. 2021) or denoising (Pang et al. 2021; Zhu et al. 2025), which typically assume globally distributed degradation, shadow degradation is inherently spatially localized and non-uniform. This unique nature presents both opportunities and challenges: while non-shadow regions offer useful guidance for recovery, the corrections needed for shadow regions often differ greatly from those for well-lit regions, making uniform processing strategies ineffective.
@@
Figure 1: Illustrates four different Mamba scanning strategies on the SRD dataset: (a) VMamba (Liu et al. 2024b), sequential scanning with PSNR = 34.70; (b) LocalMamba (Huang et al. 2024), global scanning across windows with PSNR = 29.25; (c) ShadowMamba (Zhu, Chow, and Chuah 2024), a Mamba mechanism designed for shadow removal with PSNR = 33.63; (d) Ours, dynamically determines the scanning starting point based on the shadow mask and uses a breadth-first search algorithm to select the next patch to scan with PSNR = 35.13.
@@
Existing deep learning-based methods have made considerable progress by leveraging convolutional neural networks (CNNs) (Krizhevsky, Sutskever, and Hinton 2017) and, more recently, transformer architectures (Vaswani et al. 2017). CNN-based approaches (Chen et al. 2021; Li et al. 2023; Qu et al. 2017) often exploit local contextual cues but lack the capacity to capture long-range dependencies, which are crucial for understanding global scene structure and effectively transferring illumination and texture information across regions. Transformer-based models (Gillioz et al. 2020; He et al. 2021), in contrast, excel at capturing global interactions but often suffer from high computational cost, and their attention mechanisms may lack adaptability to the spatial heterogeneity introduced by shadows. More critically, both paradigms often apply a shared processing strategy across the entire image, overlooking the unique transformation requirements of shadowed versus non-shadowed regions. Recent attempts to address this limitation have explored multi-scale processing (Ding et al. 2018; Lu et al. 2023) or guided attention mechanisms (Fukui et al. 2019); however, these typically suffer from suboptimal fusion schemes or lack the ability to adaptively propagate context based on transformation similarity.
@@
Recently, the State Space Model (SSM) architecture Mamba has attracted widespread attention for simultaneously offering long-range modeling capability and linear computational complexity. As shown in Figure 1, for vision applications, researchers have explored various Mamba scanning strategies, including rigid row/column scans(Liu et al. 2024b), fixed local-window scans(Huang et al. 2024), and region-grouped scans(Zhu, Chow, and Chuah 2024). However, these methods struggle with issues like disrupting spatial adjacency or failing to capture global dependencies effectively. As a result, a more tailored framework that integrates transformation-aware adaptation and efficient context propagation remains highly desirable.
@@
Based on this insight, we propose D2-Mamba, a novel Mamba-based network architecture specifically designed for shadow removal. Our design centers around two key innovations: (1) a Dual-Scale Fusion Mamba Block (DFMB) that fuses original and low-resolution features to enhance multi-scale representation and reduce boundary artifacts; and (2) a Dual-Path Mamba Group (DPMG) module that captures global structure using horizontal scanning while incorporating a mask-aware adaptive scanning strategy to preserve fine details and structural continuity across shadow and non-shadow regions. This strategy ensures that patches from the same region (shadow or non-shadow) are more closely grouped in the Mamba sequence, enhancing detail restoration. These two scanning mechanisms complement each other, improving the robustness of the entire module. As a result, the model can effectively capture shadow boundary information, reduce artifacts, and significantly enhance shadow removal quality. Extensive experiments on multiple benchmarks show that our method surpasses state-of-the-art models in PSNR, SSIM, and RMSE, while also delivering notable gains in efficiency and visual fidelity, validating its effectiveness for real-world shadow removal.
@@
Related Work
Shadow Removal. Early deep learning-based shadow removal methods mainly relied on CNNs to map shadowed images to shadow-free counterparts. For example, DeshadowNet (Qu et al. 2017) uses three jointly trained subnetworks to integrate semantic information for improved restoration. T-CGAN (Wang, Li, and Yang 2018) leverages stacked CGANs (Mirza and Osindero 2014) to generate both a mask and a shadow-free image, enhancing performance through mutual task improvement. Mask-ShadowGAN (Hu et al. 2019) incorporates masks into the GAN (Goodfellow et al. 2020) framework, refining shadow region localization. Despite progress, CNN-based methods struggle with limited receptive fields, failing to capture long-range dependencies, particularly in complex illumination conditions. Transformer-based architectures (Vaswani et al. 2017) alleviates these issues by capturing long-range dependencies. HomoFormer (Xiao et al. 2024a) improves local self-attention adaptability by homogenizing shadow distributions, while ShadowFormer (Guo et al. 2023a) uses global attention to enhance boundary restoration. However, the high computational complexity of Transformers hinders their practical use in high-resolution tasks.
State Space Models (SSMs). SSMs have gained attention due to low computational complexity and strong ability to model long-range dependencies. The Structured State Space Sequence Model (S4) (Gu, Goel, and Ré 2021) introduced a linear-complexity hidden state update, boosting adoption in language and vision tasks (Sarem et al. 2024; Zhang et al. 2025; Nasiri-Sarvi et al. 2024). Mamba further improved this with parallel computation and a refined update scheme.
SSMs have recently been applied to low-level vision tasks such as image super-resolution (Lei, Zhang, and Cao 2024; Qiao et al. 2024; Xiao et al. 2024b) and denoising (Lu et al. 2024; Zhu et al. 2025), showing significant improvements in computational efficiency and reconstruction quality. MambaIR (Vasluianu et al. 2024) introduced a selective SSM approach for image restoration, enhancing reconstruction quality. LocalMamba (Huang et al. 2024) utilized a local window scanning strategy to improve local interactions while retaining global context. Hi-mamba (Qiao et al. 2024) proposed a hierarchical Mamba structure for better multi-scale information fusion and feature representation.
For shadow removal, general Mamba models often fail to account for the unique characteristics of shadow images, disrupting semantic relationships at boundaries. ShadowMamba (Zhu, Chow, and Chuah 2024) introduced a selective scanning strategy for region-based partitioning, but its rigid partitioning limits adaptability. To address these issues, we enhance cross-layer information fusion and introduce a mask-aware scanning mechanism for better structural continuity and region modeling.
@@
Dual-Path Mamba Group
@@
DeshadowNet (Qu et al. 2017) - - 11.78 - - 4.84 - - 6.64
DHAN (Chen et al. 2021) 33.67 0.978 8.94 34.79 0.979 4.80 30.51 0.949 5.67
BMNet (Zhu et al. 2022) 35.05 0.981 6.61 36.02 0.982 3.61 31.69 0.956 4.46
SGNet (Wan et al. 2022) 33.76 0.979 7.45 36.48 0.984 3.05 31.39 0.960 4.23
ShadowFormer-L (Guo et al. 2023a) 36.91 0.989 5.90 36.22 0.989 3.44 32.90 0.958 4.04
ShadowDiffusion (Guo et al. 2023b) 38.72 0.987 4.98 37.78 0.985 3.44 34.73 0.970 3.63
HomoFormer (Xiao et al. 2024a) 38.81 0.987 4.25 39.45 0.988 2.85 35.37 0.972 3.33
Liu et al. (Liu et al. 2024a) 36.51 0.983 5.49 37.71 0.986 3.00 33.48 0.967 3.66
ShadowMamba (Zhu, Chow, and Chuah 2024) 37.29 0.986 5.81 37.52 0.985 3.13 33.63 0.965 3.87
ShadowGAN-Former (Hu et al. 2025) 35.24 0.981 6.578 36.12 0.981 3.54 31.73 0.956 4.39
ShadowMaskFormer (Li et al. 2025) 37.71 0.988 5.55 38.23 0.984 2.98 34.43 0.968 3.64
MSRDNet (Huang et al. 2025) 35.43 0.984 5.98 36.23 0.989 3.38 32.17 0.965 4.09
OmniSR (Xu et al. 2025) - - - - - - 34.56 0.977 -
@@
Table 1: The quantitative results of shadow removal on SRD dataset. The best and the second results are in bold and underlined.
@@
DHAN (Chen et al. 2021) 32.92 0.988 9.6 27.15 0.971 7.4 25.66 0.956 7.8
SGNet (Wan et al. 2022) 36.79 0.990 5.9 35.57 0.977 2.9 32.45 0.962 3.4
BMNet (Zhu et al. 2022) 37.87 0.991 5.8 37.51 0.985 2.4 33.98 0.972 3.0
ShadowFormer-L (Guo et al. 2023a) 39.67 - 5.2 38.82 - 2.3 35.46 - 2.8
ShadowDiffusion (Guo et al. 2023b) 39.8 - 4.9 38.90 - 2.3 35.72 - 2.7
HomoFormer (Xiao et al. 2024a) 39.49 0.991 4.9 38.64 0.982 2.3 35.31 0.969 2.6
Liu et al. (Liu et al. 2024a) 38.04 0.990 5.7 39.15 0.984 2.3 34.96 0.968 2.9
ShadowMamba (Zhu, Chow, and Chuah 2024) - - 5.8 - - 2.3 - - 2.8
MSRDNet (Huang et al. 2025) 38.93 0.991 5.5 38.49 0.985 2.4 34.94 0.972 2.9
OmniSR (Xu et al. 2025) - - - - - - 34.20 0.973 -
@@
Table 2: The quantitative results of shadow removal on the ISTD+ dataset.
@@
Dual-Scale Fusion Mamba Block
@@
Previous studies (Ren et al. 2024; Guo et al. 2023a) demonstrate that cross-scale information is crucial for achieving consistent brightness and color in shadow removal. To further incorporate cross-scale information into the Mamba framework, we draw inspiration from the position-aligned cross-scale scanning strategy proposed by MambaCSR (Ren et al. 2024). Building on this idea, we design an improved cross-scale feature fusion strategy to leverage multi-scale features and enhance the restoration capability effectively.
@@
Experiments
Experimental Setups
Datasets. We conduct experiments on three widely used shadow removal benchmarks: SRD (Qu et al. 2017), ISTD (Wang, Li, and Yang 2018), and ISTD+(Le and Samaras 2019). The SRD dataset contains 3,088 shadow/shadow-free image pairs. Since it lacks ground-truth shadow masks, we use predicted masks from DHAN (Chen et al. 2021) as substitutes. We follow the standard split with 2,680 pairs for training and 408 for testing. The ISTD dataset provides 1,870 triplets, each consisting of a shadow image, its mask, and a corresponding shadow-free image, divided into 1,330 training and 540 testing samples. ISTD+ is an improved version of ISTD, in which illumination inconsistencies between shadow and shadow-free images have been corrected through post-processing techniques.
Implementation Details. Our method is optimized using the Adam optimizer (Kingma and Ba 2014) and implemented in the PyTorch framework. All experiments were conducted on a single NVIDIA RTX 4090 GPU. The initial learning rate for training is set to $2\times 10^{-4}$, with momentum parameters $\beta_{1}$ and $\beta_{2}$ set to 0.9 and 0.999, respectively, and a batch size of 4. To optimize the training process, a cosine annealing strategy is employed to gradually reduce the learning rate, ultimately converging to $1\times 10^{-6}$.

### v3

@@ I Introduction
Shadows are common in natural images and often lead to visually distracting artifacts that adversely affect both human perception [3, 31] and the performance of high-level vision tasks [8, 46, 33]. The goal of shadow removal is to restore clean, shadow-free images from inputs that are only partially degraded. Unlike other low-level restoration tasks, such as deblurring [51, 2] or denoising [34, 50], which typically assume globally distributed degradation, shadow degradation is inherently spatially localized and non-uniform. This unique nature presents both opportunities and challenges: while non-shadow regions offer useful guidance for recovery, the corrections needed for shadow regions often differ greatly from those for well-lit regions, making uniform processing strategies ineffective.
@@
Figure 1: Illustrates four different Mamba scanning strategies on the SRD dataset: (a) VMamba [26], sequential scanning with PSNR = 34.70; (b) LocalMamba [17], global scanning across windows with PSNR = 29.25; (c) ShadowMamba [48], a Mamba mechanism designed for shadow removal with PSNR = 33.63; (d) Ours, dynamically determines the scanning starting point based on the shadow mask and uses a breadth-first search algorithm to select the next patch to scan with PSNR = 35.13.
@@
Existing deep learning-based methods have made considerable progress by leveraging convolutional neural networks (CNNs) [20] and, more recently, transformer architectures [40]. CNN-based approaches [1, 24, 36] often exploit local contextual cues but lack the capacity to capture long-range dependencies, which are crucial for understanding global scene structure and effectively transferring illumination and texture information across regions. Transformer-based models [7, 13], in contrast, excel at capturing global interactions but often suffer from high computational cost, and their attention mechanisms may lack adaptability to the spatial heterogeneity introduced by shadows. More critically, both paradigms often apply a shared processing strategy across the entire image, overlooking the unique transformation requirements of shadowed versus non-shadowed regions. Recent attempts to address this limitation have explored multi-scale processing [4, 28] or guided attention mechanisms [6]. However, these typically suffer from suboptimal fusion schemes or fail to propagate context based on transformation similarity.
@@
Recently, the State Space Model (SSM) architecture Mamba has attracted widespread attention for simultaneously offering long-range modeling capability and linear computational complexity. As shown in Figure 1, for vision applications, researchers have explored various Mamba scanning strategies, including rigid row/column scans[26], fixed local-window scans[17], and region-grouped scans[48]. However, these methods struggle with issues like disrupting spatial adjacency or failing to capture global dependencies effectively. As a result, a more tailored framework that integrates transformation-aware adaptation and efficient context propagation remains highly desirable.
@@
Based on this insight, we propose D2-Mamba, a novel Mamba-based network architecture specifically designed for shadow removal. Our design centers around two key innovations: (1) a Dual-Scale Fusion Mamba Block (DFMB) that fuses original and low-resolution features to enhance multi-scale representation and reduce boundary artifacts; and (2) a Dual-Path Mamba Group (DPMG) module that captures global structure using horizontal scanning while incorporating a mask-aware adaptive scanning strategy to preserve fine details and structural continuity across shadow and non-shadow regions. This strategy ensures that patches from the same region (shadow or non-shadow) are more closely grouped in the Mamba sequence, enhancing detail restoration. These two scanning mechanisms complement each other, improving the robustness of the entire module. As a result, the model can effectively capture shadow boundary information, reduce artifacts, and significantly enhance shadow removal quality. Extensive experiments on multiple benchmarks show that our method outperforms state-of-the-art models, while also delivering notable gains in efficiency and visual fidelity, validating its effectiveness for real-world shadow removal.
@@
II Related Work
II-A Shadow Removal
Early deep learning-based shadow removal methods mainly relied on CNNs to map shadowed images to shadow-free counterparts. For example, DeshadowNet [36] uses three jointly trained subnetworks to integrate semantic information for improved restoration. T-CGAN [42] leverages stacked CGANs [30] to generate both a mask and a shadow-free image, enhancing performance through mutual task improvement. Mask-ShadowGAN [15] incorporates masks into the GAN [9] framework, refining shadow region localization. Despite progress, CNN-based methods struggle with limited receptive fields, failing to capture long-range dependencies, particularly in complex illumination conditions. Transformer-based architectures [40] alleviates these issues by capturing long-range dependencies. HomoFormer [43] improves local self-attention adaptability by homogenizing shadow distributions, while ShadowFormer [11] uses global attention to enhance boundary restoration. However, the high computational complexity of Transformers hinders their practical use in high-resolution tasks.
II-B State Space Models in Computer Vision
SSMs have gained attention due to low computational complexity and strong ability to model long-range dependencies. The Structured State Space Sequence Model (S4) [10] introduced a linear-complexity hidden state update, boosting adoption in language and vision tasks [38, 47, 32]. Mamba further improved this with parallel computation and a refined update scheme. SSMs have recently been applied to low-level vision tasks such as image super-resolution [23, 35, 44] and denoising [29, 50], showing significant improvements in computational efficiency and reconstruction quality. MambaIR [39] introduced a selective SSM approach for image restoration, enhancing reconstruction quality. LocalMamba [17] utilized a local window scanning strategy to improve local interactions while retaining global context. Hi-mamba [35] proposed a hierarchical Mamba structure for better multi-scale information fusion and feature representation.
For shadow removal, general Mamba models often fail to account for the unique characteristics of shadow images, disrupting semantic relationships at boundaries. ShadowMamba [48] introduced a selective scanning strategy for region-based partitioning, but its rigid partitioning limits adaptability. To address these issues, we enhance cross-layer information fusion and introduce a mask-aware scanning mechanism for better structural continuity and region modeling.
@@ III-B Overall Architecture
III-C Dual-Path Mamba Group
@@
TABLE I: The quantitative results of shadow removal on SRD dataset. The best and the second results are in bold and underlined.
@@
DeshadowNet [36] - - 11.78 - - 4.84 - - 6.64
DHAN [1] 33.67 0.978 8.94 34.79 0.979 4.80 30.51 0.949 5.67
BMNet [49] 35.05 0.981 6.61 36.02 0.982 3.61 31.69 0.956 4.46
SGNet [41] 33.76 0.979 7.45 36.48 0.984 3.05 31.39 0.960 4.23
ShadowFormer-L [11] 36.91 0.989 5.90 36.22 0.989 3.44 32.90 0.958 4.04
ShadowDiffusion [12] 38.72 0.987 4.98 37.78 0.985 3.44 34.73 0.970 3.63
HomoFormer [43] 38.81 0.987 4.25 39.45 0.988 2.85 35.37 0.972 3.33
Liu et al. [27] 36.51 0.983 5.49 37.71 0.986 3.00 33.48 0.967 3.66
ShadowMamba [48] 37.29 0.986 5.81 37.52 0.985 3.13 33.63 0.965 3.87
ShadowGAN-Former [14] 35.24 0.981 6.578 36.12 0.981 3.54 31.73 0.956 4.39
ShadowMaskFormer [25] 37.71 0.988 5.55 38.23 0.984 2.98 34.43 0.968 3.64
MSRDNet [18] 35.43 0.984 5.98 36.23 0.989 3.38 32.17 0.965 4.09
OmniSR [45] - - - - - - 34.56 0.977 -
@@
TABLE II: The quantitative results of shadow removal on the ISTD+ dataset.
@@
DHAN [1] 32.92 0.988 9.6 27.15 0.971 7.4 25.66 0.956 7.8
SGNet [41] 36.79 0.990 5.9 35.57 0.977 2.9 32.45 0.962 3.4
BMNet [49] 37.87 0.991 5.8 37.51 0.985 2.4 33.98 0.972 3.0
ShadowFormer-L [11] 39.67 - 5.2 38.82 - 2.3 35.46 - 2.8
ShadowDiffusion [12] 39.8 - 4.9 38.90 - 2.3 35.72 - 2.7
HomoFormer [43] 39.49 0.991 4.9 38.64 0.982 2.3 35.31 0.969 2.6
Liu et al. [27] 38.04 0.990 5.7 39.15 0.984 2.3 34.96 0.968 2.9
ShadowMamba [48] - - 5.8 - - 2.3 - - 2.8
MSRDNet [18] 38.93 0.991 5.5 38.49 0.985 2.4 34.94 0.972 2.9
OmniSR [45] - - - - - - 34.20 0.973 -
@@
III-D Dual-Scale Fusion Mamba Block
TABLE III: The quantitative results of shadow removal using our models and recent methods on the ISTD dataset.
Method Shadow Region (S) Non-Shadow Region (NS) All Image (ALL)
DHAN [1] 34.65 0.983 8.26 29.81 0.937 5.56 28.15 0.913 6.37
BMNet [49] 35.61 0.988 7.60 32.80 0.976 4.59 30.28 0.959 5.02
ShadowFormer-S [11] 37.99 0.990 6.16 33.89 0.980 3.90 31.81 0.967 4.27
ShadowDiffusion [12] 40.15 0.994 4.13 33.70 0.977 4.14 32.33 0.969 4.12
HomoFormer [43] 39.07 0.991 5.64 33.45 0.979 4.10 31.82 0.968 4.33
ShadowMamba [48] 38.01 0.990 5.95 33.78 0.974 3.90 31.79 0.960 4.23
ShadowMaskFormer [25] - - 6.08 - - 3.86 - - 4.23
ShadowGAN-Former [14] 36.64 0.990 6.82 33.60 0.980 4.00 30.97 0.967 4.46
MSRDNet [18] 37.36 0.989 6.38 33.38 0.980 4.09 31.16 0.965 4.48
OmniSR [45] - - - - - - 31.56 0.965 -
D2-Mamba (Ours) 39.14 0.992 5.61 34.10 0.981 3.81 32.18 0.969 4.06
@@ III-D Dual-Scale Fusion Mamba Block
Previous studies [37, 11] demonstrate that cross-scale information is crucial for achieving consistent brightness and color in shadow removal. To further incorporate cross-scale information into the Mamba framework, we draw inspiration from the position-aligned cross-scale scanning strategy proposed by MambaCSR [37]. Building on this idea, we design an improved cross-scale feature fusion strategy to leverage multi-scale features and enhance the restoration capability effectively.
@@
IV Experiments
IV-A Experimental Setups
Datasets. We conduct experiments on three widely used shadow removal benchmarks: SRD [36], ISTD [42], and ISTD+[21]. The SRD dataset contains 3,088 shadow/shadow-free image pairs. Since it lacks ground-truth shadow masks, we use predicted masks from DHAN [1] as substitutes. We follow the standard split with 2,680 pairs for training and 408 for testing. The ISTD dataset provides 1,870 triplets, each consisting of a shadow image, its mask, and a corresponding shadow-free image, divided into 1,330 training and 540 testing samples. ISTD+ is an improved version of ISTD, in which illumination inconsistencies between shadow and shadow-free images have been corrected through post-processing techniques.
Implementation Details. Our method is optimized using the Adam optimizer [19] and implemented in the PyTorch framework. All experiments were conducted on a single NVIDIA RTX 4090 GPU. The initial learning rate for training is set to $2\times 10^{-4}$, with momentum parameters $\beta_{1}$ and $\beta_{2}$ set to 0.9 and 0.999, respectively, and a batch size of 4. To optimize the training process, a cosine annealing strategy is employed to gradually reduce the learning rate, ultimately converging to $1\times 10^{-6}$.
Evaluation Metrics. To enable quantitative comparisons with existing methods, we follow previous studies [16, 5, 22, 49, 41] and adopt RMSE in the LAB color space to evaluate shadow removal performance. This metric measures the difference between the predicted shadow-free images and the ground truth, where a lower RMSE indicates better restoration quality. In addition, to assess performance in the RGB color space, we also employ PSNR and SSIM as evaluation metrics. All images used for evaluation are uniformly resized to 256×256 to ensure consistency across comparisons.
@@ IV-A Experimental Setups
Figure 5: Shadow removal results of the proposed D2-Mamba on the SBU Shadow Dataset. Each pair shows the input shadow image and the corresponding shadow-free result produced by our model. The results demonstrate the model’s ability to handle complex shadow patterns under diverse real-world indoor lighting conditions.
Figure 6: Visual examples of the shadow removal results on the ISTD dataset.
IV-B Comparison with State-of-the-Art
We compare the proposed model with popular or state-of-the-art shadow removal methods. Results of all comparison methods are either obtained from their original publications or reproduced using official implementations
Quantitative evaluation. Table I, Table II and Table III present the quantitative evaluation results over SRD, ISTD+ and ISTD, respectively. Across all three datasets, our method consistently outperforms or matches existing approaches in multiple evaluation metrics, demonstrating its effectiveness and robustness in shadow removal.
On the SRD dataset, which includes complex scenes and textures, diffusion-based methods perform well due to their generative capacity for modeling intricate details. However, these methods suffer from long inference times and limited scalability for high-resolution inputs. In contrast, our Mamba-based regression model, equipped with accurate cross-scale fusion and similarity-aware grouping, achieves competitive or superior performance with significantly fewer inference steps, making it more efficient and practical for real-world deployment. On the ISTD+ dataset, our method leads in all metrics, with notable improvements in both the shadow and non-shadow regions. We achieve the highest PSNR, SSIM, and the lowest RMSE, outpacing state-of-the-art methods like ShadowDiffusion and ShadowMamba. This demonstrates the robustness of our approach in handling varied shadow types and complex lighting conditions. On the ISTD dataset, we maintain top-tier performance, outperforming or matching existing methods like ShadowFormer and ShadowDiffusion. Our approach excels in shadow regions (S) and overall image quality (ALL), with a balanced trade-off between high PSNR and SSIM scores while keeping the RMSE low. Our D2-Mamba model offers superior shadow removal quality across three datasets, balancing efficiency with high performance.
Qualitative evaluation. To further validate the effectiveness of our proposed D2-Mamba framework, we provide qualitative comparisons on the SRD, ISTD+, and ISTD datasets, which cover diverse scenes with varying shadow shapes, densities, and lighting conditions. These datasets allow for a comprehensive evaluation of the model’s robustness.
Figures 34, and 6 clearly demonstrate that our method consistently outperforms other approaches in terms of shadow boundary smoothness and artifact removal. In the SRD and ISTD+ datasets, while some methods produce shadow-recovered regions that appear noticeably brighter or darker than their surroundings (e.g., the first and third examples in Figure 3), or struggle with boundary artifacts (e.g., the second and fourth examples), our approach effectively removes residual shadows, restoring natural luminance and color for improved global consistency and visual realism. Additionally, challenging scenes with high-saturation areas (Figure 4) often lead to color inconsistencies in competing methods, whereas our method maintains color integrity across both shadow and non-shadow regions.On the ISTD dataset, D2-Mamba demonstrates strong performance in preserving texture details and structural consistency. Compared methods often produce visible residual artifacts, particularly in complex shadow regions, where shadows fail to blend seamlessly with the background or suffer from significant color distortions. In contrast, our method achieves more accurate and visually coherent shadow removal, producing natural-looking results with fewer artifacts. The shadow areas are smoothly integrated into the surrounding context, preserving the global semantic structure while accurately modeling local shadow variations.
IV-C Real-World Shadow Removal Performance
TABLE IV: Ablation study of the proposed D2-Mamba on the SRD dataset.